### v1

@@ 6.2 Numerical scheme
As a proxy for the infinite domain considered in our linear analysis, we choose periodic boundary conditions on a square domain, $[0,L]^{2}$. The initial conditions are taken as a homogeneous state $(C,P)$, where $C$ and $P$ are typically of order $\mathcal{O}(10^{-1})$, with a perturbation of random Gaussian noise with zero mean and standard deviation $10^{-3}$. For the numerical integration, we first discretise space into a grid of $100\times 100$ mesh points, and then integrate the $100^{2}$ ODEs in time using SciPy’s BDF1-5 implicit variable-order scheme. The nonlocal term is computed using a fast Fourier transform convolution method. See Jewell [21] for further details of the numerical scheme.
@@ 6.3 Results
Fig. 3 shows examples of systems in which parallel chase-and-run dynamics prevent pattern formation, while the same parameters with chiral chase-and-run lead to pattern formation. This is predicted by the dispersion relations (a, b) and is confirmed by simulations, where patterns only emerge under chiral dynamics (c-h). For the chiral simulations, the chasing and running angles are set to $\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$, although any sufficiently large angles would enable pattern formation. As evidenced in Fig. 3, this is a general phenomenon holding across different model types (direct and gradient sensing), kernel functions (tophat and exponential), as well as different combinations of self-attraction and self-repulsion in the chasers and runners. These results support our linear analysis prediction that chirality promotes pattern formation in chase-and-run systems.
@@
Another notable result from these simulations is that every stationary late-time pattern in Figs.  3, 4, and 5 consists of only a single aggregate per species. This is consistent with the results of Potts and Painter [47] and Salmaniw et al. [49], who showed that, in the absence of growth dynamics, steady-state solutions in a similar nonlocal model with a single species in 1D should have only a single peak. More generally, there is evidence that mass-conserving systems often evolve toward a single aggregate steady-state, even in local models [20, 38]. In our models, the final aggregates of the two species can either be separated (Fig. 3 c, e, f, g and Fig. 5) or mixed (Fig. 3 d, h, and Fig. 4 f). In the next section, we explore whether this outcome can be predicted by linear theory.
@@ 7 Linear analysis of separation or mixing of species
If $\Lambda_{\rho\rho}-Dk^{2}>\Lambda_{cc}-k^{2}$, then by Eq. (21), $c$ and $\rho$ are in phase.
If $\Lambda_{\rho\rho}-Dk^{2}<\Lambda_{cc}-k^{2}$, then by Eq. (22), $c$ and $\rho$ are out of phase.
@@
Figure 6: Simulations of chase-and-run with self-attracting chasers and runners. In a and c, stronger chaser self-attraction ($\mu_{cc}>\mu_{\rho\rho}$) leads to species separation at early times, as predicted by linear theory. This persists in a but not in c, where cross-species chasing is much higher than running ($|\mu_{c\rho}|\gg|\mu_{\rho c}|$), leading to mixing at $t\approx 3$ and a mixed state at late times. Conversely, in b and d, stronger runner self-attraction ($\mu_{\rho\rho}>\mu_{cc}$) causes mixing at early times. This persists in b but not in d, where cross-species running is much higher than chasing ($|\mu_{\rho c}|\gg|\mu_{c\rho}|$), leading to separation at $t\approx 1.5$ and a separated state at late times. The colourmap from Fig. 3i is used throughout, with colours rescaled only for $t=0.2$ heatmaps to display amplitudes of order $\mathcal{O}(10^{-3})$. All simulations use the direct sensing model, although gradient sensing simulations show the same qualitative behaviour. In all cases, $\alpha_{c\rho}=-20^{\circ}$, $\alpha_{c\rho}=70^{\circ}$; other parameters are given in Table 1 of the Appendix. Videos of these simulations are available at Jewell [21].
@@ 8.1 Chase-and-run
Figure 7: Late-time stationary patterns of simulations of two species with zero cross-interactions and parallel attractive self-interactions, for a range of initial densities, $P$ and $C$. The two species share all the same parameters and are essentially identical. As we might expect, at early times, the distributions of the two species did not look obviously correlated. However, despite interacting only through volume-filling, the two species can ultimately mix, leading to an in phase pattern at late times for a, c, and d. For each set of parameters, we ran simulations with 5 different realisations of the Gaussian noise initial conditions. We observed the same qualitative behaviour for every realisation when $P=C=0.1$ and $P=C=0.4$ (used in a and d, respectively). However, for $P=C=0.25$ (b and c), 2 realisations lead to separated spots (b) and 3 to a mixed stripe (c). The colourmap from Fig. 3i is used throughout. All simulations use the direct sensing model, although gradient sensing simulations exhibit the same behaviour. These systems are nonchiral, with $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=\alpha_{\rho c}=0$; other parameters are given in Table 1 of the Appendix. Videos of these simulations are available at Jewell [21].
@@ 8.2 Volume-filling can drive mixing
Another arguably counter-intuitive phenomenon regarding separation and mixing is observed when considering two species that have zero cross-interactions ($\mu_{\rho c}=\mu_{c\rho}=0$). In such cases, the only influence that one species can have on the other is through volume-filling effects, which are present through the terms $p_{c}(c,\rho)=p_{\rho}(c,\rho)=1-c-\rho$ in Equations (5) and (6). The linear theory in Section 7 predicts that having zero cross-interaction would lead to the location of one species being uncorrelated with the location of the other. We might naively guess that volume-filling, which is a purely nonlinear effect, could introduce bias towards species separation. Instead, we see simulations in which the two species become perfectly in phase and mixed in the late-time pattern, as shown in Fig. 7.
@@ 9.1 Population-level chase-and-run
We have only observed population-level chase-and-run behaviour emerging from systems with a Turing-wave bifurcation, that is, when $\text{Re}[\lambda(k)]>0$ and $\text{Im}[\lambda(k)]\neq 0$, $k>0$. However, the converse does not hold: in Fig. 8d, the pattern becomes stationary due to the large angle $\alpha_{\rho c}=80^{\circ}$, and although the imaginary component of the growth rate is reduced compared to the non-chiral case, it is still non-zero. A sufficiently larger angle would indeed eliminate this imaginary component, resulting in a standard Turing bifurcation, and likely a stationary pattern. However, 8d shows that the transition from population-level chase-and-run to stationary pattern can occur at smaller angles than this bifurcation point. This highlights a nonlinear influence of chirality on patterning.
Figure 8: Simulations with population chase-and-run with chirality. When cross-species running has a shorter interaction range than cross-species chasing ($\xi_{\rho c}<\xi_{c\rho}$), chasers form an aggregate that continually pursues an aggregate of runners [41]. Here, $\xi_{cc}=\xi_{\rho\rho}=\xi_{c\rho}=2$, whilst $\,\xi_{\rho c}=1$. a Without chirality ($\alpha_{\rho c}=0$) this pulse travels in a straight line. b Introducing chirality ($\alpha_{\rho c}=40^{\circ}$) induces population chase-and-run along a curved path. c Larger angles can cause the runner aggregate to orbit the chasers, which move less. d Even larger angles lead to chasers ‘catching’ runners, yielding a stationary mixed state. Arrows in the first panel of a-c roughly indicate the direction of movement. The colourmap from Fig. 3i is used. All simulations shown use the direct sensing model, although simulations with gradient sensing yield the same qualitative behaviour. In all cases, $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=0^{\circ}$; other parameters are given in Table 1 of the Appendix. Videos of these simulations are available at Jewell [21].
Figure 9: Two simulations (a,b) exhibiting travelling holes. In both cases, a single runner spot initially forms, encircled by both species and separated by a thin low-density gap in both populations (see first two panels of a). This gap then merges into a single hole that moves across the domain. After some transient movement: in the non-chiral case (a), the hole moves in a straight line; with chirality (b), it follows a curved trajectory. In both cases, this motion persist into late times. There is a higher density of chasers at the front of the hole (in the direction of motion), and a higher density of runners at the back. The colourmap from Fig. 3i is used throughout. In both cases, the initial density is near carry capacity ($P=C=0.48$), and $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=0^{\circ}$; other parameters are given in Table 1 of the Appendix. Videos of these simulations are available at Jewell [21].
@@ 10 Discussion
Appendix: Parameter values
Figure Model Type Kernel $C$ $P$ $D$ all $\xi$ $\mu_{cc}$ $\mu_{c\rho}$ $\mu_{\rho c}$ $\mu_{\rho\rho}$
3c direct tophat 0.25 0.25 4 1 50 50 -50 50
3d direct tophat 0.25 0.25 1 1 -50 50 -50 50
3e direct tophat 0.25 0.25 1 1 50 50 -50 -50
3f gradient exponential 0.25 0.25 4 0.5 40 40 -40 40
3g gradient tophat 0.25 0.25 4 1 20 20 -20 20
3h gradient tophat 0.25 0.25 1 1 -20 20 -20 20
4 i & 5 direct tophat 0.25 0.25 1 1 20 20 -20 50
4 ii direct tophat 0.25 0.25 1 1 50 20 -20 20
4 iii gradient tophat 0.25 0.25 1 1 8 8 -8 20
4 iv gradient tophat 0.25 0.25 1 1 20 8 -8 8
4 v direct exponential 0.25 0.25 1 0.5 20 20 -20 50
4 vi direct exponential 0.25 0.25 1 0.5 50 20 -20 20
6 a direct exponential 0.1 0.1 1 0.4 300 40 -120 150
6 b direct exponential 0.1 0.1 1 0.4 150 40 -120 300
6 c direct exponential 0.1 0.1 1 0.4 300 400 -30 150
6 d direct exponential 0.1 0.1 1 0.4 150 20 -500 300
7 a direct exponential 0.1 0.1 1 0.4 100 0 0 100
7 b & c direct exponential 0.25 0.25 1 0.4 100 0 0 100
7 d direct exponential 0.4 0.4 1 0.4 100 0 0 100
8 a-d direct exponential 0.05 0.05 1 see caption of 8 100 100 -100 100
9 direct exponential 0.48 0.48 1 0.4 600 400 -30 300

### v2

@@ 6.2 Numerical scheme
As a proxy for the infinite domain considered in our linear analysis, we choose periodic boundary conditions on a square domain, $0,L]^{2}$. The initial conditions are taken as a homogeneous state $(C,P)$, where $C$ and $P$ are typically of order $\mathcal{O}(10^{-1})$, with a perturbation of random Gaussian noise with zero mean and standard deviation $10^{-3}$. For the numerical integration, we first discretise space into a grid of $100\times 100$ mesh points, and then integrate the $100^{2}$ ODEs in time using SciPy’s BDF1-5 implicit variable-order scheme. The nonlocal term is computed using a fast Fourier transform convolution method. See Jewell [21] for further details of the numerical scheme.
@@ 6.3 Results
Fig. 4 shows examples of systems in which parallel chase-and-run dynamics prevent pattern formation, while the same parameters with chiral chase-and-run lead to pattern formation. This is predicted by the dispersion relations (a, b) and is confirmed by simulations, where patterns only emerge under chiral dynamics (c-h). For the chiral simulations, the chasing and running angles are set to $\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$, although any sufficiently large angles would enable pattern formation. As evidenced in Figs. 4 and 3, this is a general phenomenon holding across different model types (direct and gradient sensing), kernel functions (tophat and exponential), as well as different combinations of self-attraction and self-repulsion in the chasers and runners. These results support our linear analysis prediction that chirality promotes pattern formation in chase-and-run systems.
@@
Another notable result from these simulations is that every stationary late-time pattern in Figs.  4, 5, and 6 consists of only a single aggregate per species. This is consistent with the results of Potts and Painter [47] and Salmaniw et al. [49], who showed that, in the absence of growth dynamics, steady-state solutions in a similar nonlocal model with a single species in 1D should have only a single peak. More generally, there is evidence that mass-conserving systems often evolve toward a single aggregate steady-state, even in local models [20, 38]. In our models, the final aggregates of the two species can either be separated (Fig. 4 c, e, f, g and Fig. 6) or mixed (Fig. 4 d, h, and Fig. 5 f). In the next section, we explore whether this outcome can be predicted by linear theory.
@@ 7 Linear analysis of separation or mixing of species
If $\Lambda_{\rho\rho}-Dk^{2}>\Lambda_{cc}-k^{2}$, then by Equation (21), $c$ and $\rho$ are in phase.
If $\Lambda_{\rho\rho}-Dk^{2}<\Lambda_{cc}-k^{2}$, then by Equation (22), $c$ and $\rho$ are out of phase.
@@
Figure 7: Simulations of chase-and-run with self-attracting chasers and runners. In a and c, stronger chaser self-attraction ($\mu_{cc}>\mu_{\rho\rho}$) leads to species separation at early times, as predicted by linear theory. This persists in a but not in c, where cross-species chasing is much higher than running ($|\mu_{c\rho}|\gg|\mu_{\rho c}|$), leading to mixing that starts to appear around $t=2.4$ and is maintained into late times. Conversely, in b and d, stronger runner self-attraction ($\mu_{\rho\rho}>\mu_{cc}$) causes mixing at early times. This persists in b but not in d, where cross-species running is much higher than chasing ($|\mu_{\rho c}|\gg|\mu_{c\rho}|$), leading to separation at around $t=1.2$ and a separated state at late times. The colourmap from Fig. 4i is used throughout, with colours rescaled only for $t=0.2$ heatmaps to display amplitudes of order $\mathcal{O}(10^{-3})$. All simulations use the direct sensing model, although gradient sensing simulations show the same qualitative behaviour. In all cases, $\alpha_{c\rho}=-20^{\circ}$, $\alpha_{c\rho}=70^{\circ}$; other parameters are given in Table 1 of Appendix A. Videos of these simulations are available at Jewell [21].
@@ 8.1 Chase-and-run
Figure 8: Late-time stationary patterns of simulations of two species with zero cross-interactions and parallel attractive self-interactions, for a range of initial densities, $P$ and $C$. The two species share all the same parameters and are essentially identical. As we might expect, at early times, the distributions of the two species did not look obviously correlated. However, despite interacting only through volume-filling, the two species can ultimately mix, leading to an in phase pattern at late times for a, c, and d. For each set of parameters, we ran simulations with 5 different realisations of the Gaussian noise initial conditions. We observed the same qualitative behaviour for every realisation when $P=C=0.1$ and $P=C=0.4$ (used in a and d, respectively). However, for $P=C=0.25$ (b and c), 2 realisations lead to separated spots (b) and 3 to a mixed stripe (c). The colourmap from Fig. 4i is used throughout. All simulations use the direct sensing model, although gradient sensing simulations exhibit the same behaviour. These systems are nonchiral, with $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=\alpha_{\rho c}=0$; other parameters are given in Table 1 of Appendix A. Videos of these simulations are available at Jewell [21].
@@ 8.2 Volume-filling can drive mixing
Another arguably counter-intuitive phenomenon regarding separation and mixing is observed when considering two species that have zero cross-interactions ($\mu_{\rho c}=\mu_{c\rho}=0$). In such cases, the only influence that one species can have on the other is through volume-filling effects, which are present through the terms $\phi_{c}(c,\rho)=\phi_{\rho}(c,\rho)=1-c-\rho$ in Equations (5) and (6). The linear theory in Section 7 predicts that having zero cross-interaction would lead to the location of one species being uncorrelated with the location of the other. We might naively guess that volume-filling, which is a purely nonlinear effect, could introduce bias towards species separation. Instead, we see simulations in which the two species become perfectly in phase and mixed in the late-time pattern, as shown in Fig. 8.
@@ 9.1 Population-level chase-and-run
We have only observed population-level chase-and-run behaviour emerging from systems with a Turing-wave bifurcation, that is, when $\text{Re}[\lambda(k)]>0$ and $\text{Im}[\lambda(k)]\neq 0$, $k>0$. However, the converse does not hold: in Fig. 9d, the pattern becomes stationary due to the large angle $\alpha_{\rho c}=80^{\circ}$, and although the imaginary component of the growth rate is reduced compared to the non-chiral case, it is still non-zero. A sufficiently larger angle would indeed eliminate this imaginary component, resulting in a standard Turing bifurcation, and likely a stationary pattern. However, 9d shows that the transition from population-level chase-and-run to stationary pattern can occur at smaller angles than this bifurcation point. This highlights a nonlinear influence of chirality on patterning.
Figure 9: Simulations with population chase-and-run with chirality. When cross-species running has a shorter interaction range than cross-species chasing ($\xi_{\rho c}<\xi_{c\rho}$), chasers form an aggregate that continually pursues an aggregate of runners [41]. Here, $\xi_{cc}=\xi_{\rho\rho}=\xi_{c\rho}=2$, whilst $\xi_{\rho c}=1$. a Without chirality ($\alpha_{\rho c}=0$) this pulse travels in a straight line. b Introducing chirality ($\alpha_{\rho c}=40^{\circ}$) induces population chase-and-run along a curved path. c Larger angles can cause the runner aggregate to orbit the chasers, which move less. d Even larger angles lead to chasers ‘catching’ runners, yielding a stationary mixed state. Arrows in the first panel of a-c roughly indicate the direction of movement. The colourmap from Fig. 4i is used. All simulations shown use the direct sensing model, although simulations with gradient sensing yield the same qualitative behaviour. In all cases, $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=0^{\circ}$; other parameters are given in Table 1 of Appendix A. Videos of these simulations are available at Jewell [21].
Figure 10: Two simulations (a,b) exhibiting travelling holes. In both cases, a single runner spot initially forms, encircled by both species and separated by a thin low-density gap in both populations (see first two panels of a). This gap then merges into a single hole that moves across the domain. After some transient movement: in the non-chiral case (a), the hole moves in a straight line; with chirality (b), it follows a curved trajectory. In both cases, this motion persist into late times. There is a higher density of chasers at the front of the hole (in the direction of motion), and a higher density of runners at the back. The colourmap from Fig. 4i is used throughout. In both cases, the initial density is near carry capacity ($P=C=0.48$), and $\alpha_{cc}=\alpha_{\rho\rho}=\alpha_{c\rho}=0^{\circ}$; other parameters are given in Table 1 of Appendix A. Videos of these simulations are available at Jewell [21].
@@ 10 Discussion
Appendix A Parameter values
Figure Model Type Kernel $C$ $P$ $D$ all $\xi$ $\mu_{cc}$ $\mu_{c\rho}$ $\mu_{\rho c}$ $\mu_{\rho\rho}$ L
3 direct tophat 0.25 0.25 (0,5) 1 50 50 -50 50 8
4c direct tophat 0.25 0.25 4 1 50 50 -50 50 8
4d direct tophat 0.25 0.25 1 1 -50 50 -50 50 8
4e direct tophat 0.25 0.25 1 1 50 50 -50 -50 8
4f gradient exponential 0.25 0.25 4 0.5 40 40 -40 40 8
4g gradient tophat 0.25 0.25 4 1 20 20 -20 20 8
4h gradient tophat 0.25 0.25 1 1 -20 20 -20 20 8
5 i & 6 direct tophat 0.25 0.25 1 1 20 20 -20 50 8
5 ii direct tophat 0.25 0.25 1 1 50 20 -20 20 8
5 iii gradient tophat 0.25 0.25 1 1 8 8 -8 20 8
5 iv gradient tophat 0.25 0.25 1 1 20 8 -8 8 8
5 v direct exponential 0.25 0.25 1 0.5 20 20 -20 50 8
5 vi direct exponential 0.25 0.25 1 0.5 50 20 -20 20 8
7 a direct exponential 0.1 0.1 1 0.4 300 40 -120 150 3
7 b direct exponential 0.1 0.1 1 0.4 150 40 -120 300 3
7 c direct exponential 0.1 0.1 1 0.4 300 400 -30 150 3
7 d direct exponential 0.1 0.1 1 0.4 150 20 -500 300 3
8 a direct exponential 0.1 0.1 1 0.4 100 0 0 100 4
8 b & c direct exponential 0.25 0.25 1 0.4 100 0 0 100 4
8 d direct exponential 0.4 0.4 1 0.4 100 0 0 100 4
9 a-d direct exponential 0.05 0.05 1 see caption 100 100 -100 100 10
10 direct exponential 0.48 0.48 1 0.4 600 400 -30 300 3
@@ Appendix A Parameter values
Appendix B Hankel transform derivation
Here we show the derivation of the Hankel transform in the dispersion relation, illustrating the steps between the first and second line of Equation (8). The first line of this equation is
$\begin{split}\lambda(\boldsymbol{k})&=-k^{2}-P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi^{2}}\,ik\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\,e^{i\boldsymbol{k}\cdot\boldsymbol{s}}\mathop{}\!\mathrm{d}^{2}\boldsymbol{s},\end{split}$ (23)
where, as a reminder, $s\equiv|\boldsymbol{s}|$ and $k\equiv|\boldsymbol{k}|$, with $\boldsymbol{\hat{s}}\equiv\frac{\boldsymbol{s}}{s}$ and $\boldsymbol{\hat{k}}\equiv\frac{\boldsymbol{k}}{k}$.
To evaluate the integral in Equation (23), we convert to polar coordinates in $\boldsymbol{s}$ with new variables $(s,\theta_{s})$, where $\theta_{s}$ is chosen without loss to be the angle between $\boldsymbol{s}$ and $\boldsymbol{k}$. This means that $\boldsymbol{\hat{s}}=\text{cos}(\theta_{s})\boldsymbol{\hat{k}}+\text{sin}(\theta_{s})\boldsymbol{\hat{k}_{\perp}}$, where $\boldsymbol{\hat{k}_{\perp}}$ is the unit vector perpendicular to $\boldsymbol{\hat{k}}$. The integral then becomes
$\begin{split}&\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\,e^{i\boldsymbol{k}\cdot\boldsymbol{s}}\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}=\int_{0}^{\infty}\int_{0}^{2\pi}\left(\text{cos}(\theta_{s})\boldsymbol{\hat{k}}+\text{sin}(\theta_{s})\boldsymbol{\hat{k}_{\perp}}\right)\,\Omega\left(\frac{s}{\xi}\right)\,e^{iks\text{cos}(\theta_{s})}s\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}\theta_{s}\\ &=\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\left[\boldsymbol{\hat{k}}\int_{0}^{2\pi}\text{cos}(\theta_{s})e^{iks\text{cos}(\theta_{s})}\mathop{}\!\mathrm{d}\theta_{s}+\boldsymbol{\hat{k}_{\perp}}\int_{0}^{2\pi}\text{sin}(\theta_{s})e^{iks\text{cos}(\theta_{s})}\mathop{}\!\mathrm{d}\theta_{s}\right]\mathop{}\!\mathrm{d}s\,.\end{split}$ (24)
Using the change of variables $\theta_{s}=\theta^{\prime}-\pi$, we can see that the second angular integral evaluates to zero:
[…]
where the second equality arises from the integrand being an anti-symmetric function integrated across symmetric limits. Equation (24) thus simplifies to
$\begin{split}\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\,e^{i\boldsymbol{k}\cdot\boldsymbol{s}}\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&=\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\left[\boldsymbol{\hat{k}}\int_{0}^{2\pi}\text{cos}(\theta_{s})e^{iks\text{cos}(\theta_{s})}\mathop{}\!\mathrm{d}\theta_{s}\right]\mathop{}\!\mathrm{d}s\\ &=\boldsymbol{\hat{k}}\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,\left[-i\frac{\partial}{\partial(ks)}\int_{0}^{2\pi}e^{iks\text{cos}(\theta_{s})}\mathop{}\!\mathrm{d}\theta_{s}\right]\mathop{}\!\mathrm{d}s\\ &=\boldsymbol{\hat{k}}\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,\left[-i\frac{\partial}{\partial(ks)}2\pi J_{0}(ks)\right]\mathop{}\!\mathrm{d}s\\ &=\boldsymbol{\hat{k}}\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,\left[i\,2\pi J_{1}(ks)\right]\mathop{}\!\mathrm{d}s,\end{split}$ (26)
where in the third line we used the standard integral form of the $0^{\text{th}}$ order Bessel function of the first kind, $J_{0}$, and in the fourth line we used the fact that $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}x}J_{0}(x)=-J_{1}(x)$. Substituting this result back into the dispersion relation yields
$\lambda(\boldsymbol{k})=-k^{2}+2\pi\,P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi^{2}}\,k\,(\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}})\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,J_{1}(ks)\mathop{}\!\mathrm{d}s,$ (27)
which is the second line of Equation (8), as required. For further discussion of the role of Bessel functions and Hankel transforms in the linear stability analysis of these models, see Section 3 of Jewell et al. [22].
Appendix C Kernel assumptions
In order for $\Lambda_{uv}(k)$ to share the same sign as $\mu_{uv}$ for values of $k$ that are ‘relevant for pattern formation’ – typically those corresponding to the largest positive values of $\lambda(k)$, we require that: (a) $\phi_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big\rvert_{V}>0$, which follows from our core model assumptions, and (b) the Hankel transforms of the interaction kernels are positive at these key values of $k$.
Condition (b) automatically holds for any kernel whose Hankel transform is always positive. Cho et al. [6] outline some sufficiency conditions for this. For the direct sensing model, the primary condition is that $\sqrt{s}\Omega\left(\frac{s}{\xi}\right)$ is non-negative and decreasing. One way of satisfying this condition is by making a ‘splattergun’ assumption: as we increase distance, $s$, then in 2D space, interactions are spread out over a circle of increasingly large circumference $2\pi s$, and so we might expect the interaction kernel to decay at least as fast as $\frac{1}{s}$, analogously to how light intensity and some physical forces decay as $\frac{1}{s^{2}}$ in 3D space. For the gradient sensing model, in order for the $0^{\text{th}}$ order Hankel transform to always be positive, there is the same condition that $\sqrt{s}\tilde{\Omega}\left(\frac{s}{\xi}\right)$ is non-negative and decreasing, as well as an additional requirement that $\sqrt{s}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\tilde{\Omega}\left(\frac{s}{\xi}\right)$ is non-negative and decreasing. Biologically speaking, all of the above conditions correspond to interactions that do not change direction with distance, and get sufficiently weaker as distance increases. Cho et al. [6] also stipulate that certain integrals of the kernel remain finite and that the kernel vanishes at infinity, both of which are naturally satisfied in biological contexts.
The above conditions are sufficient but not necessary, and many commonly studied kernels that do not satisfy them still have Hankel transforms that are always positive. In direct sensing models, for instance, both the exponential kernel and the Gaussian-derivative kernel, $\Omega\left(\frac{s}{\xi}\right)\propto\frac{s}{\xi}e^{-0.5(\frac{s}{\xi})^{2}}$, used by Painter et al. [40], Jewell et al. [22], Salmaniw et al. [49], have positive $1^{\text{st}}$ order Hankel transforms. Similarly, for gradient sensing models, the exponential kernel and the Gaussian kernel, $\tilde{\Omega}\left(\frac{s}{\xi}\right)\propto e^{-0.5(\frac{s}{\xi})^{2}}$, used by Potts [46], Fagan et al. [11], Wang and Salmaniw [61], both have positive $0^{\text{th}}$ order Hankel transforms. Even in the case of the tophat kernel, where the Hankel transform can be negative for some values of $k$, the values of $k$ for which this occurs often do not correspond to those with the highest values of $\lambda(k)$, as discussed by Jewell et al. [22]. Essentially, the negative values of the transform have smaller magnitude than the positive values, and also occur at higher values of $k$, where the diffusion term, $k^{2}$, is larger, further reducing $\lambda$.